# Prompt-Hacking: The New p-Hacking?

### v1

@@ Abstract.
As Large Language Models (LLMs) become increasingly embedded in empirical research workflows, their use as analytical tools raises pressing concerns for scientific integrity. This opinion paper draws a parallel between “prompt-hacking”, the strategic tweaking of prompts to elicit desirable outputs from LLMs, and the well-documented practice of “p-hacking” in statistical analysis. We argue that the inherent biases, non-determinism, and opacity of LLMs make them unsuitable for data analysis tasks demanding rigor, impartiality, and reproducibility. We emphasize how researchers may inadvertently, or even deliberately, adjust prompts to confirm hypotheses while undermining research validity. We advocate for a critical view of using LLMs in research, transparent prompt documentation, and clear standards for when LLM use is appropriate. We discuss how LLMs can replace traditional analytical methods, whereas we recommend that LLMs should only be used with caution, oversight, and justification.
Large Language Models, p-Hacking, HCI Research, Research Methods, Research Validity
copyright: ccjournalyear: 2018isbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Human computer interaction (HCI)
@@ 3. Manipulating Research Results with p-Hacking
In empirical research, “p-hacking” emerges as a substantial threat to scientific integrity. P-hacking occurs when researchers tune experimental data or statistical analysis to achieve a significant p-value, a statistical measure often used to confirm or reject hypotheses (Simmons et al., 2011). Such tuning can involve selectively reporting variables, increasing sample sizes, or testing hypotheses that were changed after obtaining the results, which skew results towards significance, potentially misleading interpretations and conclusions. The consequences impact fields that rely on empirical evidence by eroding trust in findings that intensify the replication crisis, even resulting in documenting popular p-hacking strategies (Stefan and Schönbrodt, 2023). As LLMs gain prominence as research analysis tools, the potential for similar manipulation through prompt-hacking grows. We are concerned that LLMs may not be trustworthy empirical data analysis tools.
@@ 4. Prompt-Hacking: p-Hacking with LLMs
We state that “prompt-hacking” closely resembles “p-hacking,” a problematic practice in data analysis where researchers tune variables, data, and statistical tests to achieve significant p-values. Prompt-hacking phenomena were introduced recently (Kosch and Feger, 2024; Morris, 2024), and much like p-hacking, prompt-hacking may subconsciously encourage selective data manipulation. For example, researchers could keep modifying prompts to obtain outputs that support desired conclusions. Morris stated in a related opinion article, “Prompting is a poor user interface for LLMs, which should be phased out as quickly as possible” (Morris, 2024). Researchers, especially non-LLM experts, may be unaware of how to prompt and how slight distinctions between prompting and natural-language interaction may create different research results. Unlike traditional research methods, LLM outputs vary dramatically depending on prompt phrasing and style. This variability in pseudo-natural language poses a challenge for reproducibility. Each prompt, even if only slightly altered, can yield different outputs, making it nearly impossible to replicate findings reliably. As Morris noted, the lack of transparency in documenting prompt variations, validation processes, and final prompt selection biases can damage the scientific integrity of empirical studies. Prior studies explored using LLMs for data analysis (Tabone and de Winter, 2023) and even for simulating human subject experiments (Aher et al., 2023). Yet, the prompting space is infinite, and subtle semantic or syntactic prompt changes may provide different research results. Morris highlighted that failing to report the number and history of unsuccessful prompts along with any distinguishing features of successful ones, neglecting to test whether slight prompt variations affect outcomes, and not verifying prompt consistency across different models, model versions, or repeated uses of the same model all represent significant oversights for research replicability when using LLMs.
@@ 5. Are LLMs Appropriate for Data Analysis?
While structured guidelines can mitigate some risks, researchers must be cautious of over-reliance on LLMs for tasks requiring impartiality. These models are different from human judgment and traditional qualitative or quantitative analysis. It is important to understand that even with improved transparency and documentation, the fundamental limitations of LLMs mean they should be used sparingly and primarily as a supplement to human analysis, not a substitute. LLMs are here to stay, and it is likely, or even already the case, that researchers rely on LLMs as a research tool (Schmidt et al., 2024; Tabone and de Winter, 2023). We urge future research directions to advocate for careful LLM use. Although novel scientific insights can mitigate the risks of prompt-hacking, researchers must remain vigilant about the fundamental limitations of LLMs. Unlike p-hacking, where the misuse of statistical techniques can be uncovered through reproducible means, using LLMs as data analysts inherently introduces biases and inaccuracies, even when following guidelines, due to their non-deterministic output. We propose that researchers adopt a cautious ’just don’t do it’ mindset in cases where LLM use introduces unnecessary risks or could replace established rigorous methods. Only in limited and justified scenarios, where the benefits of LLMs outweigh their risks, should LLMs be considered tools for analysis.
@@
Preregistration and Documentation of Prompts: Based on the prompt stability process, researchers should preregister prompts and experimental protocols to ensure transparency. This includes documenting the sequence of prompts and their post-modifications deviating from the preregistration, helping prevent selective disclosure of prompts that favor specific hypotheses. While preregistration and documentation help reduce PARKing, the core issue is deciding whether LLMs should be used. Researchers must resist the temptation to repeatedly adjust prompts to align results with hypotheses. Instead, they should critically evaluate whether the task requires LLM. The answer will often be that LLMs are unnecessary and potentially harmful. All these recommendations affect researchers, publication outlets, funding agencies, and research infrastructure providers. Infrastructure providers, including Zenodo and the Center for Open Science, must extend their features to capture preregistered prompts and metadata on target LLMs and their precise versions.
@@ 6. Moving Towards Ethical and Reliable Use of LLMs in Research
Whether to trust LLMs as impartial data analysts demands a clear and cautious stance: no, they should only be trusted with significant oversight. While their utility in accelerating specific research processes is undeniable, their inherent biases and variability show the need for a restrained approach and more research in this area. Researchers must prioritize the integrity of the scientific process above convenience, actively questioning the role and limitations of LLMs in empirical research. While the comparison to p-hacking highlights similarities in the risks of manipulation, it is essential to stress a key distinction: the outputs of LLMs are fundamentally shaped by their design and training, making them less objective than statistical tools. Unlike p-hacking, which often involves misused but inherently neutral techniques, prompt-hacking exploits tools that are not impartial by design. As such, even the “correct” use of LLMs in analysis cannot guarantee validity, demanding caution and critical oversight. The central question is not how to use LLMs responsibly but whether they should be used. The answer for most data analysis tasks is clear: avoid LLMs unless their use is essential and justifiable. The scientific community must resist the temptation to normalize LLM-based analysis and instead uphold the rigor and integrity of traditional methods. The safest course for most data analysis tasks is straightforward: just don’t do it.

### v2

@@ Abstract.
As Large Language Models (LLMs) become increasingly embedded in empirical research workflows, their use as analytical tools for quantitative or qualitative data raises pressing concerns for scientific integrity. This opinion paper draws a parallel between “prompt-hacking”, the strategic tweaking of prompts to elicit desirable outputs from LLMs, and the well-documented practice of “p-hacking” in statistical analysis. We argue that the inherent biases, non-determinism, and opacity of LLMs make them unsuitable for data analysis tasks demanding rigor, impartiality, and reproducibility. We emphasize how researchers may inadvertently, or even deliberately, adjust prompts to confirm hypotheses while undermining research validity. We advocate for a critical view of using LLMs in research, transparent prompt documentation, and clear standards for when LLM use is appropriate. We discuss how LLMs can replace traditional analytical methods, whereas we recommend that LLMs should only be used with caution, oversight, and justification.
Large Language Models, Reproducibility, Prompt-Hacking, p-Hacking, Data Analysis
copyright: rightsretainedjournalyear: 2025doi: 10.1145/3744911ccs: Human-centered computing Human computer interaction (HCI)
@@ 3. Manipulating Research Results with p-Hacking
In empirical research, p-hacking emerges as a substantial threat to scientific integrity. P-hacking occurs when researchers tune experimental data or statistical analysis to achieve a significant p-value, a statistical measure often used to confirm or reject hypotheses (Simmons et al., 2011). Such tuning can involve selectively reporting variables, increasing sample sizes, or testing hypotheses that were changed after obtaining the results, which skew results towards significance, potentially misleading interpretations and conclusions. The consequences impact fields that rely on empirical evidence by eroding trust in findings that intensify the replication crisis, even resulting in documenting popular p-hacking strategies (Stefan and Schönbrodt, 2023). As LLMs gain prominence as research analysis tools, the potential for similar manipulation through prompt-hacking grows. We are concerned that LLMs may not be trustworthy empirical data analysis tools.
@@ 4. Prompt-Hacking: p-Hacking with LLMs
We state that “prompt-hacking” closely resembles “p-hacking,” a problematic practice in data analysis where researchers tune variables, data, and statistical tests to achieve significant p-values. Prompt-hacking phenomena were introduced recently (Kosch and Feger, 2024; Morris, 2024), and much like p-hacking, prompt-hacking may subconsciously encourage selective data manipulation. For example, researchers could keep modifying prompts to obtain outputs that support desired conclusions. Morris stated in a related opinion article, “Prompting is a poor user interface for LLMs, which should be phased out as quickly as possible” (Morris, 2024). Researchers, especially non-LLM experts, may be unaware of how to prompt and how slight distinctions between prompting and natural-language interaction may create different research results. Unlike traditional research methods, LLM outputs vary dramatically depending on prompt phrasing and style. This variability in pseudo-natural language poses a challenge for reproducibility. Each prompt, even if only slightly altered, can yield different outputs, making it impossible to replicate findings reliably. As Morris noted, the lack of transparency in documenting prompt variations, validation processes, and final prompt selection biases can damage the scientific integrity of empirical studies. Prior studies explored using LLMs for data analysis (Tabone and de Winter, 2023) and even for simulating human subject experiments (Aher et al., 2023). Yet, the prompting space is infinite, and subtle semantic or syntactic prompt changes may provide different research results. Morris highlighted that failing to report the number and history of unsuccessful prompts along with any distinguishing features of successful ones, neglecting to test whether slight prompt variations affect outcomes, and not verifying prompt consistency across different models, model versions, or repeated uses of the same model all represent significant oversights for research replicability when using LLMs.
@@ 5. Are LLMs Appropriate for Data Analysis?
While structured guidelines can mitigate some risks, researchers must be cautious of over-reliance on LLMs for tasks requiring impartiality. These models are different from human judgment and traditional qualitative or quantitative analysis. It is important to understand that even with improved transparency and documentation, the fundamental limitations of LLMs mean they should be used sparingly and primarily as a supplement to human analysis, not a substitute. LLMs are here to stay, and it is likely, or even already the case, that researchers rely on LLMs as a research tool (Schmidt et al., 2024; Tabone and de Winter, 2023). We urge future research directions to advocate for careful LLM use. Although novel scientific insights can mitigate the risks of prompt-hacking, researchers must remain vigilant about the fundamental limitations of LLMs. Unlike p-hacking, where the misuse of statistical techniques can be uncovered through reproducible means, using LLMs as data analysts inherently introduces biases and inaccuracies, even when following guidelines, due to their non-deterministic output. We propose that researchers adopt a cautious mindset in cases where LLM use introduces unnecessary risks or could replace established rigorous methods. Only in limited and justified scenarios, where the benefits of LLMs outweigh their risks, should they be considered as tools for analysis.
@@
Preregistration and Documentation of Prompts: Based on the prompt stability process, researchers should preregister prompts and experimental protocols to ensure transparency. This includes documenting the sequence of prompts and their post-modifications deviating from the preregistration, helping prevent selective disclosure of prompts that favor specific hypotheses. While preregistration and documentation help reduce PARKing, the core issue is deciding whether LLMs should be used. Researchers must resist the temptation to repeatedly adjust prompts to align results with hypotheses. Instead, they should critically evaluate whether the task requires LLMs. The answer will often be that LLMs are unnecessary and potentially harmful. All these recommendations affect researchers, publication outlets, funding agencies, and research infrastructure providers. Infrastructure providers, including Zenodo and the Center for Open Science, must extend their features to capture preregistered prompts and metadata on target LLMs and their precise versions.
@@ 6. Moving Towards Ethical and Reliable Use of LLMs in Research
Whether to trust LLMs as impartial data analysts demands a clear and cautious stance: no, they should only be trusted with significant oversight. While their utility in accelerating specific research processes is undeniable, their inherent biases and variability show the need for a restrained approach and more research in this area. Researchers must prioritize the integrity of the scientific process above convenience, actively questioning the role and limitations of LLMs in empirical research. While the comparison to p-hacking highlights similarities in the risks of manipulation, it is essential to stress a key distinction: the outputs of LLMs are fundamentally shaped by their design and training, making them less objective than statistical tools. Unlike p-hacking, which often involves misused but inherently neutral techniques, prompt-hacking exploits tools that are not impartial by design. As such, even the “correct” use of LLMs in analysis cannot guarantee validity, demanding caution and critical oversight. The central question is not how to use LLMs responsibly but whether they should be used. The answer for most data analysis tasks is clear: avoid LLMs unless their use is essential and justifiable. The scientific community must resist the temptation to normalize LLM-based analysis and instead uphold the rigor and integrity of traditional methods.
7. Acknowledgements
This work is supported by the German Research Foundation (DFG), CRC 1404: “FONDA: Foundations of Workflows for Large-Scale Scientific Data Analysis” (Project-ID 414984028).